# LLM-PBE: Assessing Data Privacy in Large Language Models

### v1

@@ 4.4. Practicality of PETs on Fine-tuning of LLMs
We investigate the effectiveness of scrubbing in mitigating privacy risks. Specifically, we fine-tune Llama-2 7b on the ECHR dataset for 4 epochs and use four MIA approaches (PPL, Refer, LiRA, and MIN-K) with ECHR to assess privacy leakage from the fine-tuned model. Our focus is on the impact of these techniques on privacy leakage, without considering potential overfitting. The results, presented in Table 4, indicate that scrubbing and DP can effectively reduce the MIA AUC. However, we observe that the scrubbing process significantly degrades model performance, highlighting a critical challenge in balancing privacy protection and model utility.
@@ 5. Leakage of Prompts
Leaking in-store system prompts can cause significant financial losses. In this section, we conduct a comprehensive evaluation of prompt privacy using different PLAs methods, models, and potential defenses. We focus on answering the following research questions: 1) Is prompt easily leaked using attack prompts? 2) How does the risk of prompt leakage vary across different LLMs? 3) Is it possible to protect the prompts by using defensive prompting?
@@ 5.1. Experimental Setup
Metrics. We follow (Perez and Ribeiro, 2022) to measure the extraction quality by the RapidFuzz package (Bachmann, 2021). RapidFuzz leverages the Levenshtein Distance to calculate the similarity between two strings, which is informally the minimum number of single-character edits (insertions, deletions, or substitutions) required to change one string into the other. For brevity, we call the similarity score as FuzzRate (FR). The similarity score ranges from 0 to 100 (fully matched). If each text is randomly shuffled, the score will be 83.9 on average over 300 samples from BlackFriday.
@@ 5.4. Effectiveness of Defensive Prompting
Mitigation for private-information breach. Breach of private information through the leaked prompt can be mitigated by using privacy-preserving algorithms in generating prompts (Hong et al., 2023; Tang et al., 2023; Panda et al., 2023a). This usually involves the use of private samples as in-context learning examples. DP-OPT (Hong et al., 2023) is the first end-to-end prompt tuning solution, that uses an offsite small model to generate prompts by learning from private data. DP-ICL Generation (Tang et al., 2023) utilizes in-context learning to generate insensitive samples by LLMs for specific tasks. Rather than doing training or synthesizing data, DP-ICL (Panda et al., 2023a) directly ensembles multiple subsets of private samples to generate responses. All three methods leverage DP to account and bound privacy costs.
@@ 7. Conclusions
In conclusion, our paper has thoroughly explored the data privacy risks associated with LLMs. We provide a systematic toolkit to assess the data privacy of LLMs, which can be easily adopted by LLM researchers and developers. Through a comprehensive analysis of various attack and defense methodologies, we have identified key trends and vulnerabilities in LLM privacy. Our study underscores the evolving nature of these risks and the increasing importance of developing more robust privacy-preserving mechanisms in this field. The insights gained from our research not only highlight the complexities inherent in securing LLMs but also pave the way for future advancements in this domain. By systematically documenting and analyzing the current state of LLM privacy, our work serves as a crucial reference for further exploration and innovation, aiming to balance the remarkable capabilities of these models with the imperative of protecting data privacy.
@@ Acknowledgement
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP2020-018), Infocomm Media Development Authority under its Trust Tech Funding Initiative, Singapore National Research Foundation funding #053424, ARL funding #W911NF-23-2-0137, DARPA funding #112774-19499, the National Science Foundation under grant no. 2229876 as well as no. 2212176, and more funds provided by the National Science Foundation, by the Department of Homeland Security, by IBM, by Berkeley Center for Responsible Decentralized Intelligence (RDI), and by TogetherAI. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the supporting entities.
@@ A.1. Data Extraction Attacks
Poisoning-Based Methods. Poisoning-based methods assume that the attacker can modify the training data to insert poisons with a similar pattern as the target secret, and train the model on the poisoned dataset.  Panda et al. (2023b) show that an attacker can inject poisons into a training dataset that induce the model to memorize the secret (e.g., PII) that is unknown to the attacker during training, and then easily extract this memorized secret during inference. Similarly, in (Jayaraman et al., 2022), each of the poison points is a message–response pair (i.e., Email Id, Password, Credential) that has a recurring pattern in the response part, similar to the sensitive data the attacker is trying to extract. When the LLM is trained on this message–response pair, it is likely to memorize the password pattern and associate the prefix pattern password with the actual sensitive password.
@@ A.3.1. Manually Designed Prompts
Input Obfuscation: The goal of the jailbreaking prompts in this category is to obfuscate the attack goal so that LLMs cannot detect the query as a malicious query (Wei et al., 2023). There are three main approaches to obfuscate the attack goal: encoding, splitting, and role play. 1) Encoding-based methods: attackers encode the query and provide the encoded query and description of the encoding method to LLMs. The encoding approach can be one of the existing approaches that LLMs can understand (e.g, Base64, Morse code) or a custom method where the encoding function should also be fed into the LLM (e.g., a mapping function) (Yuan et al., 2023). 2) Splitting-based methods: attackers split the attack keywords into multiple subparts so that LLMs cannot detect them (Kang et al., 2023). For example, while directly inputting “social security card” can be easily detected by LLMs, we can assign “social” to a variable A, “security” to a variable B, and “card” to a variable C. Then, we ask the LLM to combine the string A+B+C and answer it. 3) Role play-based methods: attackers ask the LLM to act as a given character in a specified scenario. A representative example is DAN (red, 2023), where the prompt asks the LLM to act as DAN, which stands for “do anything now”, and respond to user queries without any restrictions. To enhance the LLM to act in the given role, examples can be provided in the query to let the LLM know how the role would respond (Li et al., 2023a).
@@ A.3.2. Automatic Prompt Generation
Poisoning-based Training Extraction rate Pythia, GPT-2, Bert2Bert (Panda et al., 2023b; Jayaraman et al., 2022)
@@
Jailbreaking Input obfuscation Post-training Attack success rate GPT-3.5/4 (Wei et al., 2023; Yuan et al., 2023; Kang et al., 2023; red, 2023; Li et al., 2023a)
@@ Appendix B Summarization of Defense Approaches
DP in-context learning (Panda et al., 2023a)
@@
Machine unlearning Modified training (Bourtoule et al., 2021; Kumar et al., 2022)
Fine-tuning (Jang et al., 2022; Wang et al., 2023b; Warnecke et al., 2021)
@@ B.1. Differential Privacy
DP for Finetuning. Yu et al. (2021) defend against privacy leakage of finetuning data when releasing private finetuned models. Lukas et al. (2023) provide comprehensive experiments to evaluate the PII leakage when fine-tuning models on sensitive data. DP-SGD often suffers from high dimensions not only from computation overhead but also performance (Li et al., 2022). When LLMs have been very memory-intensive, clipping gradients sample-wisely in DP-SGD will severely increase the demand for larger memory consumption. There is a series of works aiming to improve the memory efficiency of DP-SGD on LLMs (Li et al., 2021; Bu et al., 2022a, 2023, b). Instead of improving the parameter complexity, Yu et al. (2023a) consider reducing the size of samples to be protected with non-private data.
DP Prompt Tuning There is increasing interest in incorporating differential privacy (DP) (Dwork, 2006) into prompt tuning for privacy protection. PromptPATE utilized a DP ensemble approach to label public data. Using these as in-context examples, they devised a discrete prompt tailored for few-shot learning on designated models (Duan et al., 2023). In parallel, DP In-Context Learning (Panda et al., 2023a) advocates for ensembling multiple in-context samples to predict classification labels. However, both works assume a set of non-private data which may not hold in practice. In the absence of public datasets, Duan et al. (2023) also show the viability of DP-SGD (Abadi et al., 2016) in the realm of soft prompt tuning.  Li et al. (2023b) propose to paraphrase prompts rendering a sample-wise notion of privacy.
@@ B.3. Machine Unlearning
Unlearning through Modified Training Machine unlearning approaches within this category require modifications to the original training process. One classic approach (Bourtoule et al., 2021; Kumar et al., 2022) is to partition the training data and train a model on each partitioned data. The ensemble of the trained models is used for prediction. Then, when some data are required to be unlearned, only the partitions that involve the deleted data are affected, and only the corresponding models need to be retained, which reduces the computation cost compared with retraining on the remained data. However, such kind of approach has not been applied to LLMs yet due to the expensive cost of retraining.
Unlearning through Fine-Tuning Machine unlearning approaches within this category fine-tune the trained model to unlearn the deleted data without modification to the original training process. There are several approaches (Jang et al., 2022; Wang et al., 2023b; Warnecke et al., 2021) on how to fine-tune the model. 1) Gradient ascent on the deleted data (Jang et al., 2022): given the deleted token sequence, the unlearning objective is to minimize the probability of correctly predicting the next token given part of the deleted token sequence. 2) Knowledge gap alignment (Wang et al., 2023b): the model is updated such that the knowledge gap between it and the model trained on the deleted data is similar to the gap of another model handling the seen and unseen data. Specifically, given a model $M_{d}$ trained on the deleted data $D_{d}$, a small set of extra data $D_{e}$ which is not in the training data and a model $M_{e}$ trained on $D_{e}$, the trained model $M_{o}$ on the original data $D_{o}$ is updated to minimize the gap between the difference of the output distributions between $M_{o}$ and $M_{e}$ on $D_{e}$ and the difference of the output distributions between the current model and $M_{d}$ on $D_{d}$. 3) Unlearning with the influence of the deleted data (Warnecke et al., 2021): considering removing features of labels, the model is updated to fill the influence of the change of data. The objective is to minimize the difference between the loss of the model on the changed data and the loss of the model on the original data. Both first-order and second-order updates are considered in the work.

### v2

@@ 4.4. Practicality of PETs on Fine-tuning of LLMs
We investigate the effectiveness of scrubbing in mitigating privacy risks. Specifically, we fine-tune Llama-2 7b on the ECHR dataset for 4 epochs and use four MIA approaches (PPL, Refer, LiRA, and MIN-K) with ECHR to assess privacy leakage from the fine-tuned model. We use FastDP to implement the DP-SGD with better memory and computation efficiency [^5] https://github.com/awslabs/fast-differential-privacy/. Our focus is on the impact of these techniques on privacy leakage, without considering potential overfitting. The results, presented in Table 5, indicate that scrubbing and DP can effectively reduce the MIA AUC. However, we observe that the scrubbing process significantly degrades model performance, highlighting a critical challenge in balancing privacy protection and model utility.
@@ 5. Leakage of Prompts
Leaking in-store system prompts can cause significant financial losses. In this section, we conduct a comprehensive evaluation of prompt privacy using different PLA methods, models, and potential defenses. We focus on answering the following research questions: 1) Is prompt easily leaked using attack prompts? 2) How does the risk of prompt leakage vary across different LLMs? 3) Is it possible to protect the prompts by using defensive prompting?
@@ 5.1. Experimental Setup
Metrics. We follow (Perez and Ribeiro, 2022) to measure the extraction quality by the RapidFuzz package (Bachmann, 2021). RapidFuzz leverages the Levenshtein Distance to calculate the similarity between two strings, which is informally the minimum number of single-character edits (insertions, deletions, or substitutions) required to change one string into the other. For brevity, we call the similarity score as FuzzRate (FR). The similarity score ranges from 0 to 100 (fully matched). If two sample include the same word set but the words are randomly shuffled, the score will be 83.9 on average over 300 samples from BlackFriday. Thus, a FR larger than 83.9 implies a good match in the word set.
@@ 5.4. Effectiveness of Defensive Prompting
Mitigation for private-information breach. Breach of private information through the leaked prompt can be mitigated by using privacy-preserving algorithms in generating prompts (Hong et al., 2024; Tang et al., 2023; Panda et al., 2023). This usually involves the use of private samples as in-context learning examples. DP-OPT (Hong et al., 2024) is the first end-to-end prompt tuning solution, that uses an offsite small model to generate prompts by learning from private data. DP-ICL Generation (Tang et al., 2023) utilizes in-context learning to generate insensitive samples by LLMs for specific tasks. Rather than doing training or synthesizing data, DP-ICL (Panda et al., 2023) directly ensembles multiple subsets of private samples to generate responses. All three methods leverage DP to account and bound privacy costs.
@@ 7. Conclusions
In conclusion, our paper has thoroughly explored the data privacy risks associated with LLMs. We provide a systematic toolkit to assess the data privacy of LLMs, which can be easily adopted by LLM researchers and developers. Through a comprehensive analysis of various attack and defense methodologies, we have identified key trends and vulnerabilities in LLM privacy. Our study underscores the evolving nature of these risks and the increasing importance of developing more robust privacy-preserving mechanisms in this field. The insights gained from our research not only highlight the complexities inherent in securing LLMs but also pave the way for future advancements in this domain.
@@ Acknowledgement
We thank Zhiqi Bu @Amazon AI for his comments during our development. This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP2020-018), Infocomm Media Development Authority under its Trust Tech Funding Initiative, Singapore National Research Foundation funding #053424, ARL funding #W911NF-23-2-0137, DARPA funding #112774-19499, the National Science Foundation under grant no. 2229876 as well as no. 2212176, and more funds provided by the National Science Foundation, by the Department of Homeland Security, by IBM, by Berkeley Center for Responsible Decentralized Intelligence (RDI), and by TogetherAI. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the supporting entities.
@@ A.1. Data Extraction Attacks
Poisoning-Based Methods. Poisoning-based methods assume that the attacker can modify the training data to insert poisons with a similar pattern as the target secret, and train the model on the poisoned dataset.  Panda et al. (2024) show that an attacker can inject poisons into a training dataset that induce the model to memorize the secret (e.g., PII) that is unknown to the attacker during training, and then easily extract this memorized secret during inference. Similarly, in (Jayaraman et al., 2022), each of the poison points is a message–response pair (i.e., Email Id, Password, Credential) that has a recurring pattern in the response part, similar to the sensitive data the attacker is trying to extract. When the LLM is trained on this message–response pair, it is likely to memorize the password pattern and associate the prefix pattern password with the actual sensitive password.
@@ A.3.1. Manually Designed Prompts
Input Obfuscation: The goal of the jailbreaking prompts in this category is to obfuscate the attack goal so that LLMs cannot detect the query as a malicious query (Wei et al., 2023). There are three main approaches to obfuscate the attack goal: encoding, splitting, and role play. 1) Encoding-based methods: attackers encode the query and provide the encoded query and description of the encoding method to LLMs. The encoding approach can be one of the existing approaches that LLMs can understand (e.g, Base64, Morse code) or a custom method where the encoding function should also be fed into the LLM (e.g., a mapping function) (Yuan et al., 2023). 2) Splitting-based methods: attackers split the attack keywords into multiple subparts so that LLMs cannot detect them (Kang et al., 2024). For example, while directly inputting “social security card” can be easily detected by LLMs, we can assign “social” to a variable A, “security” to a variable B, and “card” to a variable C. Then, we ask the LLM to combine the string A+B+C and answer it. 3) Role play-based methods: attackers ask the LLM to act as a given character in a specified scenario. A representative example is DAN (red, 2023), where the prompt asks the LLM to act as DAN, which stands for “do anything now”, and respond to user queries without any restrictions. To enhance the LLM to act in the given role, examples can be provided in the query to let the LLM know how the role would respond (Li et al., 2023a).
@@ A.3.2. Automatic Prompt Generation
Poisoning-based Training Extraction rate Pythia, GPT-2, Bert2Bert (Panda et al., 2024; Jayaraman et al., 2022)
@@
Jailbreaking Input obfuscation Post-training Attack success rate GPT-3.5/4 (Wei et al., 2023; Yuan et al., 2023; Kang et al., 2024; red, 2023; Li et al., 2023a)
@@ Appendix B Summarization of Defense Approaches
DP in-context learning (Panda et al., 2023)
@@
Machine unlearning Modified training (Bourtoule et al., 2021; Kumar et al., 2023)
Fine-tuning (Jang et al., 2023; Wang et al., 2023b; Warnecke et al., 2021)
@@ B.1. Differential Privacy
DP for Finetuning. Yu et al. (2021) defend against privacy leakage of finetuning data when releasing private finetuned models. Lukas et al. (2023) provide comprehensive experiments to evaluate the PII leakage when fine-tuning models on sensitive data. DP-SGD often suffers from high dimensions not only from computation overhead but also performance (Li et al., 2022a). When LLMs have been very memory-intensive, clipping gradients sample-wisely in DP-SGD will severely increase the demand for larger memory consumption. There is a series of works aiming to improve the memory efficiency of DP-SGD on LLMs (Li et al., 2022b; Bu et al., 2023c, e, d). Instead of improving the parameter complexity, Yu et al. (2023a) consider reducing the size of samples to be protected with non-private data.
DP Prompt Tuning There is increasing interest in incorporating differential privacy (DP) (Dwork, 2006) into prompt tuning for privacy protection. PromptPATE utilized a DP ensemble approach to label public data. Using these as in-context examples, they devised a discrete prompt tailored for few-shot learning on designated models (Duan et al., 2023). In parallel, DP In-Context Learning (Panda et al., 2023) advocates for ensembling multiple in-context samples to predict classification labels. However, both works assume a set of non-private data which may not hold in practice. In the absence of public datasets, Duan et al. (2023) also show the viability of DP-SGD (Abadi et al., 2016) in the realm of soft prompt tuning.  Li et al. (2023b) propose to paraphrase prompts rendering a sample-wise notion of privacy.
@@ B.3. Machine Unlearning
Unlearning through Modified Training Machine unlearning approaches within this category require modifications to the original training process. One classic approach (Bourtoule et al., 2021; Kumar et al., 2023) is to partition the training data and train a model on each partitioned data. The ensemble of the trained models is used for prediction. Then, when some data are required to be unlearned, only the partitions that involve the deleted data are affected, and only the corresponding models need to be retained, which reduces the computation cost compared with retraining on the remained data. However, such kind of approach has not been applied to LLMs yet due to the expensive cost of retraining.
Unlearning through Fine-Tuning Machine unlearning approaches within this category fine-tune the trained model to unlearn the deleted data without modification to the original training process. There are several approaches (Jang et al., 2023; Wang et al., 2023b; Warnecke et al., 2021) on how to fine-tune the model. 1) Gradient ascent on the deleted data (Jang et al., 2023): given the deleted token sequence, the unlearning objective is to minimize the probability of correctly predicting the next token given part of the deleted token sequence. 2) Knowledge gap alignment (Wang et al., 2023b): the model is updated such that the knowledge gap between it and the model trained on the deleted data is similar to the gap of another model handling the seen and unseen data. Specifically, given a model $M_{d}$ trained on the deleted data $D_{d}$, a small set of extra data $D_{e}$ which is not in the training data and a model $M_{e}$ trained on $D_{e}$, the trained model $M_{o}$ on the original data $D_{o}$ is updated to minimize the gap between the difference of the output distributions between $M_{o}$ and $M_{e}$ on $D_{e}$ and the difference of the output distributions between the current model and $M_{d}$ on $D_{d}$. 3) Unlearning with the influence of the deleted data (Warnecke et al., 2021): considering removing features of labels, the model is updated to fill the influence of the change of data. The objective is to minimize the difference between the loss of the model on the changed data and the loss of the model on the original data. Both first-order and second-order updates are considered in the work.